OpenStudy (anonymous):

Help: What is the solution to the following equation? x2 + 5x - 2 = 0

OpenStudy (anonymous):

x2 means x^2 ??

OpenStudy (anonymous):

\[(-5\pm \sqrt{5^2-4*1*-2})/2*1\]

OpenStudy (anonymous):

A. x = B. x = C. x = -1 D. x = 2

OpenStudy (anonymous):

\[x1=(-5+\sqrt{33})/2\] \[x1=(-5-\sqrt{33})/2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

answers are 0.351 and -2.851